Demand during leadtime for CD-ROM drives averages 80units ( normally distributed), with a standard deviation of 8 drives. Mr. Smith, the purchasing manager, is willing to accept a stockout risk of no more than 3 percent. What is the appropriate reorder point ( Z = 1.88)?
Answer #1

Given data:

Mean Demand (µ) = 80 units

Standard Deviation (σ) = 8

Maximum Stockout risk = 3 %

Z = 1.88

Safety stock = x - µ

Safety stock = z * σ = 1.88 * 8 = 15.04

Reorder point = Expected demand during lead time + Safety stock = Mean demand + Safety stock = 80 + 15.04 = 95.04